OpenStudy (anonymous):

Find the product of the binomial? (x3 - 3x)(5x2 + 2)

OpenStudy (whpalmer4):

Okay, to multiply polynomials like this, just remember that each part of the first polynomial multiplies each part of the second polynomial. Then collect like terms to simplify. For example: \[(x^2 - 2x)(x + 4) = x^2(x+4) - 2x(x+4) \]\[= x^2*x + x^2*4 - 2x*x -2x*4 \]\[= x^3+4x^2-2x^2-8x\]\[=x^3+2x^2-8x \]

OpenStudy (whpalmer4):

You are just using the distributive property of multiplication over and over...

OpenStudy (anonymous):

Can you multiply x^3 and 5x^2 together? @whpalmer4

OpenStudy (whpalmer4):

Yes, of course. \[ x^3 * 5x^2 = 5 * x^3 * x^2 = 5 * (x*x*x) * (x*x)= 5x^5\]

OpenStudy (whpalmer4):

By the properties of exponents, \[x^n*x^m = x^{n+m}\]

OpenStudy (whpalmer4):

But they have to have the same base — you can't do \[x^3*y^2 = x^5\] or anything like that. \[x^3*y^2 = x^3y^2\]